OpenStudy (anonymous):

Convert 6.40x10^-17 coulombs to elementary charges

OpenStudy (unklerhaukus):

\[6.40\times10^{-17}[\text{C}]\\ =6.40\times10^{-17}[\text{C}]\times\frac1{1.602\times10^{-19}\left[\frac{\text C}{\text e}\right]}\\ =\frac{6.40\times10^{-17}}{1.602\times10^{-19}}[\text e]\\ =\]